# Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM

### v1

@@ Abstract
In large-scale hypothesis testing, computing exact $p$-values or $e$-values is often resource-intensive, creating a need for budget-aware inferential methods. We propose a general framework for active hypothesis testing that leverages inexpensive auxiliary statistics to allocate a global computational budget. For each hypothesis, our data-adaptive procedure probabilistically decides whether to compute the exact test statistic or a transformed proxy, guaranteeing a valid $p$-value or $e$-value while satisfying the budget constraint in expectation. Theoretical guarantees are established for our constructions, showing that the procedure achieves optimality for $e$-values and for $p$-values under independence, and admissibility for $p$-values under general dependence. Empirical results from simulations and two real-world applications, including a large-scale genome-wide association study (GWAS) and a clinical prediction task leveraging large language models (LLM), demonstrate that our framework improves statistical efficiency under fixed resource limits.
@@ 1 Introduction
The $p$-value and $e$-value (Vovk and Wang,, 2021; Ren and Barber,, 2023; Ramdas and Wang,, 2024) are fundamental tools in statistical inference for quantifying evidence against a null hypothesis. While essential, their exact computation can be prohibitively expensive due to costly experimental procedures or substantial computational demands. This challenge creates a need for inferential methods that operate within a fixed budget. We propose a general framework for active hypothesis testing that addresses this problem directly. Our approach leverages inexpensive and readily available auxiliary statistics, which are derived from cheaper data sources, prior knowledge, or predictive models, to manage a global computational budget. For each hypothesis, a data-adaptive procedure probabilistically decides whether to compute the resource-intensive “gold-standard” statistic. When the exact statistic is not computed, a transformed version of the auxiliary statistic is used in its place, ensuring a valid test statistic for every hypothesis. This framework finds wide applicability across various domains, as illustrated by the following examples.
@@
    Powerful Prediction Model. Consider a setting where a large, pre-trained prediction model is available to forecast an outcome of interest (Angelopoulos et al.,, 2023; Motwani and Witten,, 2023; Zrnic and Candès,, 2024; Kluger et al.,, 2025; Ji et al.,, 2025). The exact $e$-value or $p$-value requires observing the actual outcomes, which may be costly or delayed. A proxy statistic can be rapidly computed by using the model’s predictions in place of the true outcomes. In this setting, the proxy statistic acts as an auxiliary statistic to guide whether observing the true outcomes is worth the cost or delay.
@@
    Costly vs. Noisy Measurements. In many scientific and industrial domains, a precise measurement is destructive, time-consuming, or financially expensive, while a cheaper but noisier measurement is often available from alternative sensors (Carroll et al.,, 1995; Fuller,, 2009; Grace et al.,, 2021; Dunbar et al.,, 2022). For instance, a full genetic assay is costly, but a simple biomarker measurement is not. A valid $p$-value or $e$-value can only be derived from the precise measurement, while the noisy data provides an informative, yet potentially biased, proxy. Our framework formally navigates this trade-off, using the noisy measurement as an informative guide to determine when the budget should be spent on the definitive, precise measurement.
@@
    Multi-view Learning and Complementary Signals. In many applications, data provide multiple distinct “views” of the same underlying phenomenon, where different perspectives can offer complementary information (Zhang et al.,, 2011; Sun,, 2013; Zhao et al.,, 2017). A common setting involves two complementary data views. The first view, e.g., genomic or genotype data, is expensive to collect but supports valid computation of $p$-values or $e$-values. The second view, such as routine clinical measurements or gene expression data, is inexpensive to obtain but its null distribution is unknown, making it unsuitable for direct inferential procedures. Despite this limitation, the second view can provide substantial predictive signal. Our framework synthesizes these complementary sources of information by using the cheap data view as a powerful proxy to strategically allocate the budget for computations on the resource-intensive view.
A central challenge addressed in this work is the efficient integration of two types of information: a costly but statistically valid test statistic, and an auxiliary statistic that is inexpensive to obtain but may be less reliable. Formally, we consider a setting with $N$ hypotheses, each associated with a resource-intensive test statistic that yields valid inference and a cheap auxiliary statistic that may be unreliable. Our objective is to construct a valid test statistic for every hypothesis while ensuring that the expected number of costly computations does not exceed a predetermined global budget.
@@ 1.1 Related Work
Prior work on incorporating auxiliary statistics in hypothesis testing has largely focused on improving statistical power, with limited consideration of computational budget constraints. These approaches typically leverage side information to prioritize hypotheses, and are often implemented through weighted multiple testing procedures, which can be interpreted either as re-weighting the $p$-values (Genovese et al.,, 2006; Ignatiadis et al.,, 2016; Liu et al.,, 2016; Barber and Ramdas,, 2017; Xia et al.,, 2020; Cai et al.,, 2022) or, equivalently, as adaptively adjusting the rejection thresholds (Lei and Fithian,, 2018; Zhang et al.,, 2019; Li and Barber,, 2019; Chao and Fithian,, 2021; Freestone et al.,, 2024). Although these approaches improve power, they rely on the assumption that an exact $p$-value is available for every hypothesis and therefore do not tackle the fundamental challenge of high computational or experimental cost. Nevertheless, the weights or prioritization scores generated by some of these approaches can be leveraged as auxiliary statistics within our framework to inform efficient budget allocation. Similarly, two-stage multiple testing procedures use an inexpensive screening stage to filter out unpromising hypotheses (Zehetmayer et al.,, 2005; Aoshima and Yata,, 2011). These methods, however, typically rely on a hard selection rule: hypotheses that fail the initial screening are discarded, and no formal inferential statements are made for them.
A second line of research, which inspires our approach, is active learning, where information is queried selectively to improve efficiency (Cohn et al.,, 1996; Settles,, 2009; Sener and Savarese,, 2018; Ren et al.,, 2021). However, our goal is fundamentally different. The active learning literature, including recent work on acquiring gold-standard labels for statistical inference (Zhang et al.,, 2021; Zrnic and Candès,, 2024; Cook et al.,, 2024), has primarily focused on optimizing the collection of labeled data for parameter estimation or model training. While conceptually related, these methods aim to improve the efficiency of data collection, whereas our work focuses on the dynamic decision of whether to compute a test statistic itself.
The work most closely related to ours is the recently proposed proxy computing framework of Xu et al., 2025b , which introduces an innovative approach to reducing the computational burden of hypothesis testing by probabilistically deciding whether to compute an exact statistic. Our work builds upon this foundation by introducing a key extension: we develop a flexible, global budget-constraint formulation. Whereas the original method makes independent per-hypothesis decisions, our framework guarantees adherence to a pre-specified total budget in expectation. Furthermore, we establish that our proposed construction is optimal within a broad and intuitive class of methods, thereby providing stronger theoretical guarantees.
@@ 1.2 Our Contributions
To address the aforementioned limitations, we develop a flexible and efficient framework for active hypothesis testing under a computational budget. Our approach leverages inexpensive auxiliary statistics to intelligently allocate computational resources, thereby maximizing statistical power while rigorously respecting budget constraints and maintaining full statistical validity.
@@
Our proposal offers several novel and valuable contributions. First, we establish a formal budget-constrained procedure, guaranteeing that the expected number of expensive computations does not exceed a user-specified limit. This feature makes our framework directly applicable in real-world, resource-constrained settings, providing a formal bridge between statistical objectives and practical limitations. Theoretical guarantees and extensive empirical results confirm that our method achieves tight budget adherence: the actual number of computations closely matches the specified budget, avoiding both wasteful under-utilization and over-expenditure. Second, our framework is model-free, imposing no distributional requirements on the auxiliary statistics. This property is uniquely suited for integrating unstructured information from complex black-box systems, such as LLMs, where the generative process of the auxiliary statistic is unknown or intractable. Finally, we provide rigorous theoretical guarantees for our constructions, showing that our procedure attains optimality for $e$-values and for $p$-values under independence, as well as admissibility for $p$-values under general dependence. This positions our approach as a principled and theoretically sound method, rather than a heuristic.
@@ 2.1 Problem Formulation
Consider a set of $N$ null hypotheses, $\{H_{0,i}\}_{i=1}^{N}$. For each hypothesis $H_{0,i}$, we have access to two types of statistics:
@@
Our primary objective is to generate a valid test statistic (an active $e$-value or active $p$-value) for every hypothesis $i\in\{1,\dots,N\}$, while adhering to a pre-specified global budget. We assume that each computation of an expensive statistic $X_{i}$ incurs one unit of cost. The global budget, denoted by $n_{b}$ (where typically $n_{b}\ll N$), represents an upper bound on the expected total number of costly computations.
Formally, let $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$ be an indicator variable for the decision to compute the expensive statistic for hypothesis $i$. The budget constraint is then given by:
$\mathbb{E}\left[\sum_{i=1}^{N}C_{i}\right]\leq n_{b}.$ (1)
@@ 2.2 Active $e$-value
The choice of the multiplicative form $b(E^{a})E$ is deliberate. It is designed to preserve the role of the exact $e$-value, $E$, which is typically a carefully constructed measure of evidence. This structure is intuitive and interpretable, as it simply re-scales the original evidence based on the auxiliary statistic $E^{a}$. We prefer this simple re-scaling to more complex transformations (e.g., $E^{2}$ or other nonlinear functions) that could obscure the relationship between the final statistic and the original $e$-value. Our framework also includes the active $e$-value proposed in Xu et al., 2025b as a special case, with a more detailed comparison provided in Section F of the supplement.
@@
A natural and intuitive approach to satisfying this inequality is to control each of the two terms in the sum separately. This can be achieved by partitioning the total tolerable expectation of one with a constant $\beta\in[0,1]$, bounding the first term by $\beta$ and the second by $1-\beta$. This decomposition is sufficient, since the two constraints guarantee the total expectation is bounded by $1$. The following theorem provides a complete characterization, showing that this decomposition is not just a convenient strategy but is in fact necessary for the validity of any such active $e$-value construction.
@@
For $E^{\mathrm{active}}$ as defined in Definition 1, given the control function $h(\cdot)$, the following are equivalent:
  1. 1.
    $\mathbb{E}[E^{\mathrm{active}}]\leq 1$ for all joint distributions of non-negative random variables $(E^{a},E)$ with $\mathbb{E}[E]\leq 1$.
  2. 2.
    There exists $\beta\in[0,1]$ such that: $\sup_{x\geq 0}a(x)(1-h(x))\leq\beta\text{ and }\sup_{x\geq 0}b(x)h(x)\leq 1-\beta.$
@@ 2.3 Active $p$-value
Similarly to the $e$-value case, the multiplicative form $b(P^{a})P$ is a deliberate choice. It preserves the original structure of the exact $p$-value $P$, which is often carefully constructed for high power, by simply re-scaling it (e.g., Barber and Ramdas,, 2017; Li and Barber,, 2019; Xia et al.,, 2020; Cai et al.,, 2022). Moreover, our framework encompasses the active $p$-value proposed in Xu et al., 2025b as a special case. Further discussions are provided in Section F of the supplement.
@@
The separate constraints (4) and (5) are sufficient, but not necessary, to ensure the active $p$-value satisfies the super-uniformity condition in (3). A counterexample is provided in Section E of the supplement. This contrasts with the $e$-value case in Theorem 1. While our approach imposes a stronger condition than strictly required, it furnishes a tractable framework for constructing a broad class of valid active $p$-values.
@@
In contrast, the optimal choice of $b(\cdot)$ under Condition (5) is more nuanced, as it is governed by the joint distribution of $P$ and $P^{a}$. For instance, the choice $b(q)=h(q)/(1-\beta)$, which is analogous to the optimal $e$-value construction, fails to satisfy Condition (5) under general dependence (a counterexample is provided in Section D of the supplement). This distinction motivates the need for separate constructions depending on the dependency structure, which we formalize in the following theorem.
@@
    If $P$ and $P^{a}$ are independent, the point-wise smallest $b(\cdot)$ that satisfies (5) is: $b(x)=\frac{h(x)}{1-\beta}.$
@@ 2.4 Admissibility and the Choice of Control Parameters
The active statistic constructions in (2), (6) and (7) depend on the choice of the control function $h(\cdot)$ and the hyperparameter $\beta$. A fundamental theoretical question naturally arises: does a universally optimal configuration exist? That is, can we identify a specific function $h$ and parameter $\beta$ that yield a more powerful test against all alternatives?
To answer this, we formally define statistical domination and admissibility within our framework. Intuitively, one active statistic dominates another if it is always “better”, which means yielding a larger $e$-value or smaller $p$-value regardless of the data realization.
@@
Assume $h$ is non-trivial (not identically 0 or $1$). No single $\beta\in(0,1)$ uniformly dominates all others. In fact, for any fixed $h(\cdot)$, every active statistic induced by any $\beta\in(0,1)$ is admissible.

### v2

@@ Abstract
In large-scale hypothesis testing, computing exact $p$-values or $e$-values is often resource-intensive, creating a need for budget-aware inferential methods. We propose a general framework for active hypothesis testing that leverages inexpensive auxiliary statistics to allocate a global computational budget. For each hypothesis, our data-adaptive procedure probabilistically decides whether to compute the exact test statistic or a transformed proxy, guaranteeing a valid $p$-value or $e$-value while satisfying the exact budget constraint. Theoretical guarantees are established for our constructions, showing that the procedure achieves optimality for $e$-values and for $p$-values under independence, and admissibility for $p$-values under general dependence. Empirical results from simulations and two real-world applications, including a large-scale genome-wide association study (GWAS) and a clinical prediction task leveraging large language models (LLM), demonstrate that our framework improves statistical efficiency under fixed resource limits.
@@ 1 Introduction
The $p$-value and $e$-value (Vovk and Wang, 2021; Ren and Barber, 2023; Ramdas and Wang, 2024) are fundamental tools in statistical inference for quantifying evidence against a null hypothesis. While essential, their exact computation can be prohibitively expensive due to costly experimental procedures or substantial computational demands. This challenge creates a need for inferential methods that operate within a fixed budget. We propose a general framework for active hypothesis testing that addresses this problem directly. Our approach leverages inexpensive and readily available auxiliary statistics, which are derived from cheaper data sources, prior knowledge, or predictive models, to manage a global computational budget. For each hypothesis, a data-adaptive procedure probabilistically decides whether to compute the resource-intensive “gold-standard” statistic. When the exact statistic is not computed, a transformed version of the auxiliary statistic is used in its place, ensuring a valid test statistic for every hypothesis. This framework finds wide applicability across various domains, as illustrated by the following examples.
@@
    Powerful Prediction Model. Consider a setting where a large, pre-trained prediction model is available to forecast an outcome of interest (Angelopoulos et al., 2023; Motwani and Witten, 2023; Zrnic and Candès, 2024b; Kluger et al., 2025; Ji et al., 2025). The exact $e$-value or $p$-value requires observing the actual outcomes, which may be costly or delayed. A proxy statistic can be rapidly computed by using the model’s predictions in place of the true outcomes. In this setting, the proxy statistic acts as an auxiliary statistic to guide whether observing the true outcomes is worth the cost or delay.
@@
    Costly vs. Noisy Measurements. In many scientific and industrial domains, a precise measurement is destructive, time-consuming, or financially expensive, while a cheaper but noisier measurement is often available from alternative sensors (Carroll et al., 1995; Fuller, 2009; Grace et al., 2021; Dunbar et al., 2022). For instance, a full genetic assay is costly, but a simple biomarker measurement is not. A valid $p$-value or $e$-value can only be derived from the precise measurement, while the noisy data provides an informative, yet potentially biased, proxy. Our framework formally navigates this trade-off, using the noisy measurement as an informative guide to determine when the budget should be spent on the definitive, precise measurement.
@@
    Multi-view Learning and Complementary Signals. In many applications, data provide multiple distinct “views” of the same underlying phenomenon, where different perspectives can offer complementary information (Zhang et al., 2011; Sun, 2013; Zhao et al., 2017). A common setting involves two complementary data views. The first view, e.g., genomic or genotype data, is expensive to collect but supports valid computation of $p$-values or $e$-values. The second view, such as routine clinical measurements or gene expression data, is inexpensive to obtain but its null distribution is unknown, making it unsuitable for direct inferential procedures. Despite this limitation, the second view can provide substantial predictive signal. Our framework synthesizes these complementary sources of information by using the cheap data view as a powerful proxy to strategically allocate the budget for computations on the resource-intensive view.
A central challenge addressed in this work is the efficient integration of two types of information: a costly but statistically valid test statistic, and an auxiliary statistic that is inexpensive to obtain but may be less reliable. Formally, we consider a setting with NN hypotheses, each associated with a resource-intensive test statistic that yields valid inference and a cheap auxiliary statistic that may be unreliable. Our objective is to construct a valid test statistic for every hypothesis while ensuring that the number of costly computations strictly adheres to a predetermined global budget.
@@ 1.1 Related Work
Prior work on incorporating auxiliary statistics in hypothesis testing has largely focused on improving statistical power, with limited consideration of computational budget constraints. These approaches typically leverage side information to prioritize hypotheses, and are often implemented through weighted multiple testing procedures, which can be interpreted either as re-weighting the $p$-values (Genovese et al., 2006; Ignatiadis et al., 2016; Liu et al., 2016; Barber and Ramdas, 2017; Xia et al., 2020; Cai et al., 2022) or, equivalently, as adaptively adjusting the rejection thresholds (Lei and Fithian, 2018; Zhang et al., 2019; Li and Barber, 2019; Chao and Fithian, 2021; Freestone et al., 2024). Although these approaches improve power, they rely on the assumption that an exact $p$-value is available for every hypothesis and therefore do not tackle the fundamental challenge of high computational or experimental cost. Nevertheless, the weights or prioritization scores generated by some of these approaches can be leveraged as auxiliary statistics within our framework to inform efficient budget allocation. Similarly, two-stage multiple testing procedures use an inexpensive screening stage to filter out unpromising hypotheses (Zehetmayer et al., 2005; Aoshima and Yata, 2011). These methods, however, typically rely on a hard selection rule: hypotheses that fail the initial screening are discarded, and no formal inferential statements are made for them.
A second line of research, which inspires our approach, is active learning, where information is queried selectively to improve efficiency (Cohn et al., 1996; Settles, 2009; Sener and Savarese, 2018; Ren et al., 2021). However, our goal is fundamentally different. The active learning literature, including recent work on acquiring gold-standard labels for statistical inference (Zhang et al., 2021; Zrnic and Candès, 2024a; Cook et al., 2024), has primarily focused on optimizing the collection of labeled data for parameter estimation or model training. While conceptually related, these methods aim to improve the efficiency of data collection, whereas our work focuses on the dynamic decision of whether to compute a test statistic itself.
The work most closely related to ours is the recently proposed proxy computing framework of Xu et al. (2025b), which employs probabilistic queries of exact test statistics to reduce expected computational costs. That approach, however, relies on a fixed test construction and makes query decisions independently for each hypothesis. As a result, it does not provide general optimality guarantees and the total computational cost remains stochastic. We generalize this framework by demonstrating that the construction in Xu et al. (2025b) is a special case of a broader class of valid active statistics. By characterizing this class, we establish the optimality and admissibility theory for active inference and replace the independent query mechanism with a budget-constrained global allocation.
@@ 1.2 Our Contributions
To address the aforementioned limitations, we develop a flexible and efficient framework for active hypothesis testing under a computational budget. Our approach leverages inexpensive auxiliary statistics to allocate computational resources in a way that maximizes statistical power, while strictly respecting budget constraints and maintaining statistical validity.
@@
Our work makes several contributions. First, we establish a budget-constrained procedure that guarantees the number of expensive computations exactly matches a user-specified limit on every run. Second, our framework is model-free, imposing no distributional requirements on the auxiliary statistics. This property is uniquely suited for integrating unstructured information from complex black-box systems, such as LLMs, where the generative process of the auxiliary statistic is unknown or intractable. Finally, we provide rigorous theoretical guarantees for our constructions, showing that our procedure attains optimality for $e$-values and for $p$-values under independence, as well as admissibility for $p$-values under general dependence. This positions our approach as a principled and theoretically sound method, rather than a heuristic.
@@ 2.1 Problem Formulation
Consider a set of NN null hypotheses, $\{H_{0,i}\}_{i=1}^{N}$. For each hypothesis $H_{0,i}$, we have access to two types of statistics:
@@
Our primary objective is to generate a valid test statistic (an active $e$-value or active $p$-value) for every hypothesis $i\in\{1,\dots,N\}$, while adhering to a pre-specified global budget. We assume that each computation of an expensive statistic $X_{i}$ incurs one unit of cost. The global budget, denoted by $n_{b}$ (where typically $n_{b}\ll N$), represents the total number of costly computations allowed. Formally, let $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$ be an indicator variable for the decision to compute the expensive statistic for hypothesis ii. The budget constraint is then given by:
$\sum_{i=1}^{N}C_{i}\leq n_{b}.$} (1)
@@ 2.2 Active $e$-value
The choice of the multiplicative form $b(E^{a})E$ is deliberate. It is designed to preserve the role of the exact $e$-value, $E$, which is typically a carefully constructed measure of evidence. This structure is intuitive and interpretable, as it simply re-scales the original evidence based on the auxiliary statistic $E^{a}$. We prefer this simple re-scaling to more complex transformations (e.g., $E^{2}$ or other nonlinear functions) that could obscure the relationship between the final statistic and the original $e$-value. Our framework also includes the active $e$-value proposed in Xu et al. (2025b) as a special case, with a more detailed comparison provided in Section G of the supplement.
@@
A natural and intuitive approach to satisfying this inequality is to control each of the two terms in the sum separately. This can be achieved by partitioning the total tolerable expectation of one with a constant $\beta\in[0,1]$, bounding the first term by $\beta$ and the second by $1-\beta$. This decomposition is sufficient, since the two constraints guarantee the total expectation is bounded by 11. The following theorem provides a complete characterization, showing that this decomposition is not just a convenient strategy but is in fact necessary for the validity of any such active $e$-value construction.
@@
For $E^{\mathrm{active}}$ as defined in Definition 1, given the control function $h(\cdot)$, the following two statements are equivalent: (1) $\mathbb{E}[E^{\mathrm{active}}]\leq 1$ for all joint distributions of non-negative random variables $(E^{a},E)$ with $\mathbb{E}[E]\leq 1$; (2) There exists $\beta\in[0,1]$ such that: $\sup_{x\geq 0}a(x)(1-h(x))\leq\beta\text{ and }\sup_{x\geq 0}b(x)h(x)\leq 1-\beta.$
@@ 2.3 Active $p$-value
Similarly to the $e$-value case, the multiplicative form $b(P^{a})P$ is a deliberate choice. It preserves the original structure of the exact $p$-value $P$, which is often carefully constructed for high power, by simply re-scaling it (e.g., Barber and Ramdas, 2017; Li and Barber, 2019; Xia et al., 2020; Cai et al., 2022). Moreover, our framework encompasses the active $p$-value proposed in Xu et al. (2025b) as a special case. Further discussions are provided in Section G of the supplement.
@@
The separate constraints (4) and (5) are sufficient, but not necessary, to ensure the active $p$-value satisfies the super-uniformity condition in (3). A counterexample is provided in Section F of the supplement. This contrasts with the $e$-value case in Theorem 1. While our approach imposes a stronger condition than strictly required, it furnishes a tractable framework for constructing a broad class of valid active $p$-values.
@@
In contrast, the optimal choice of $b(\cdot)$ under Condition (5) is more nuanced, as it is governed by the joint distribution of $P$ and $P^{a}$. For instance, the choice $b(q)=h(q)/(1-\beta)$, which is analogous to the optimal $e$-value construction, fails to satisfy Condition (5) under general dependence (a counterexample is provided in Section E of the supplement). This distinction motivates the need for separate constructions depending on the dependency structure, which we formalize in the following theorem.
@@
    If $P$ and $P^{a}$ are independent, the point-wise smallest $b(\cdot)$ that satisfies (5) is:
    $b(x)=\frac{h(x)}{1-\beta}.$
@@ 2.4 Admissibility and the Choice of Control Parameters
The active statistic constructions in (2), (6) and (7) depend on the choice of the control function $h(\cdot)$ and the hyperparameter $\beta$. While recent literature (Xu et al., 2025b) has introduced specific functional forms for active statistics, a rigorous theoretical evaluation of whether these or any other choices are optimal has remained absent. This naturally raises a fundamental question: does a universally optimal configuration actually exist? That is, can we identify a specific function $h$ and parameter $\beta$ that yield a strictly more powerful test against all alternatives? To answer this question and address the gap in prior work, we provide a in-depth theoretical investigation into the admissibility of active statistics. We begin by formally defining statistical domination and admissibility within our framework. Intuitively, one active statistic dominates another if it is always “better”, which means yielding a larger $e$-value or smaller $p$-value regardless of the data realization.
@@
Assume $h$ is non-trivial (not identically 0 or 11). No single $\beta\in(0,1)$ uniformly dominates all others. In fact, for any fixed $h(\cdot)$, every active statistic induced by any $\beta\in(0,1)$ is admissible.